# AC3D: Analyzing and Improving 3D Camera Control in Video Diffusion Transformers

### v2

@@ 3.5 Mitigating training data limitations
Estimating camera parameters from in-the-wild videos remains challenging, as leading methods like [190, 115, 114, 145] frequently fail when processing videos containing dynamic scene content. This limitation has resulted in camera-annotated datasets being heavily biased toward static scenes, particularly evident in RealEstate10K (RE10K) [197], which is the predominant dataset for training camera-controlled video models [149, 39, 6]. We hypothesize that models fine-tuned on such data interpret camera position information as a signal to suppress scene dynamics. This bias persists even when jointly training on unconstrained 2D video data [150], because the camera conditioning branch is only activated when camera parameters are available, which occurs exclusively for static scenes from RE10K, as static scenes remain the only reliable source for accurate camera annotation.
To address this fundamental limitation, we propose an alternative approach: rather than attempting to annotate dynamic scenes, which proved unsuccessful in our extensive preliminary research, even with state-of-the-art methods [145], we curate a collection of $20$K diverse videos featuring dynamic scenes captured by stationary cameras (see Figure 6). With stationary cameras, the camera position is inherently known, (we can assign fixed arbitrary extrinsics), allowing us to maintain active camera conditioning during training. This approach enables the camera conditioning branch to remain active during training while exposing the model to dynamic content, helping it distinguish between viewpoint conditioning and scene stillness. On top of this secondary dataset, following [150], we remove the scale ambiguity in RE10K by leveraging an off-the-shelf metric depth estimator; see Appendix G. Our experiments in Sec. 4.3 demonstrate that this straightforward yet effective data curation strategy successfully mitigates the distributional limitations of RE10K, restoring much of the lost scene dynamics, while maintaining precise camera control.
@@
Figure 6: RealEstate10k [197] videos (upper 2 rows) contain diverse camera trajectories, but are strongly biased towards static scenes. To mitigate this bias and also increase the concepts diversity, we curate $20$K videos with stationary cameras, but dynamic content (lower 2 rows). Such dataset is easy to construct, and surprisingly effective. Section 4.3 shows that integrating it into our training allows to improve visual quality on out-of-distribution prompts by $17\%$.
@@ 3.6 Miscellaneous improvements
1D temporal camera encoder
VDiT-CCprocesses videos in the autoencoder’s latent space with 4×4$\times$4 × temporal compression [167], raising the question of how to incorporate full-resolution camera parameters. While camera poses could be naively downsized (e.g., subsampled or reshaped) to match the latent resolution, this would force small DiT-XS blocks to process compressed camera information. Instead, we implement a sequence of causal 1D convolutions that transform a $F{\times}6$ sequence of Plücker coordinates for each pixel into a $(F//4){\times}32$ representation; see Appendix C.
@@
$\hat{s}(\bm{x}|\bm{y},\bm{c})$ $=(1+w_{y}+w_{c})s_{\theta}(\bm{x}|\bm{y},\bm{c})$ (1)
$-w_{y}s_{\theta}(\bm{x}|\bm{c})-w_{c}s_{\theta}(\bm{x}|\bm{y}),$
where $\hat{s}(.)$ denotes the final update direction used during synthesis, $s_{\theta}$ represents the model’s predicted update direction, $\bm{y}$ and $\bm{c}$ are text and camera conditions, and $w_{y}$ and $w_{c}$ are their respective CFG weights.
@@
Traditional ControlNet [186], used in recent camera control methods [149, 39, 6], only processes conditioning signals without accessing the main branch. Our experiments show that adding cross-attention from video tokens to camera tokens in transformer computations enables better camera representations. This modification behaves as a feedback mechanism [71] provided by the main synthesis branch to the camera processing branch.
@@
We found that applying cross-attention over the context information (text prompts, resolution, etc.) in the camera DiT-XS blocks lead to detrimental visual quality and camera steering due to harmful interference of the context embeddings with camera representations.
@@ 4 Experiments
Our base VDiT model was trained on a large-scale dataset of text-annotated images and videos. VDiT-CC is fine-tuned from VDiT on RealEstate10K [197], contains ${\approx}{65}$M video clips with per-frame camera parameters (extrinsics and intrinsics) since it is the setup used by existing methods [149, 39, 6].
@@
To assess the performance, we rely on a wide range of automatic quantitative metrics and user studies. For the metrics, we use FID [43], FVD [136], and CLIP score [42] to evaluate visual quality, and rotation and normalized translation errors [39] of ParticleSfM [190]-reconstructed trajectories to assess camera steerability. We evaluate them both on RE10K and MSR-VTT [160], since the latter allows to assess zero-shot generalization on out-of-distribution data.
  • Motion Quality (MQ): Which video has larger and more natural motion.
  • Visual Quality (VQ): Which video has a higher overall visual quality.
@@ 4.1 Baselines
We select 3 contemporary camera-control methods: MotionCtrl [149], CameraCtrl [39], and VD3D [6]. MotionCtrl and CameraCtrl use a UNet-based video diffusion backbone [37], while VD3D builds on top of FIT [93, 21] and as such, is easily extendable to our video DiT [100] setup. That’s why we re-implement it on top of our VDiT model to obtain an additional “VD3D+DiT” baseline. Besides, we also employ the unconditional base VDiT model as a reference for visual quality degradation.
Table 2: Quantitative evaluation. We evaluate all the models using camera pose and visual quality metrics based on unseen camera trajectories. We compute translation and rotation errors based on the estimated camera poses from generations using ParticleSfM [190]. We evaluate both in-distribution with RealEstate10K [197] and out-of-distribtion with MSR-VTT [160].
@@
MotionCtrl [149] 0.477 0.094 2.99 61.70 26.46 0.593 0.137 16.85 283.12 24.11
CameraCtrl [39] 0.465 0.089 2.48 55.64 26.81 0.587 0.132 12.33 201.33 25.05
VD3D (FIT) [6] 0.409 0.043 1.40 42.43 28.07 0.504 0.050 7.80 165.18 26.89
VD3D (DiT) 0.421 0.056 1.21 38.57 28.34 0.486 0.047 6.88 137.62 27.90
AC3D (ours) 0.358 0.035 1.18 36.55 28.76 0.428 0.038 5.34 110.71 28.58
@@ 4.2 Main results
We present the quantitative comparison results against the baselines in Tab. 2. One can observe that just switching from 4B-parameters pixel-space FIT [21] backbone, employed by the original VD3D approach, to our larger 11.5B-parameters latent-space DiT yielded clear gains across the most of the metrics. Next, the results demonstrate that our AC3D establishes a new state-of-the-art performance against all the baselines, and achieves ${\approx}{18}$% better visual quality than the closest competitor (VD3D DiT) and ${\approx}{25}$% more precise camera steering.  As evaluating the quality of camera motion by staring at a small selection of frames in a PDF is incredibly difficult, we instead opt for visualizing all qualitative results in the website provided within our supplementary material. Therein, we can observe that AC3D better follows pose conditioning, and achieves higher visual fidelity.  We conduct user studies against VD3D+FIT (the original model) and VD3D+DiT (our improved re-implementation on top of the bigger video transformer). The results are presented in Table 1: our AC3D outperforms them across all the qualitative aspects, achieving 90%+ overall preference score.  Finally, we encourage the reader to assess the visual quality by observing the video comparisons in our supplementary materials.
@@ 4.3 Ablations
The first ablation we conduct is dropping any camera conditioning at all, which makes the model equivalent to vanilla VDiT. It is needed to understand the visual quality and text alignment degradation. The results (Tab. 2, row w/o camera cond) show that our model loses less only ${\approx}{7}$% of the original visual fidelity on MSR-VTT(as measured by FVD), while (as expected) greatly improves on its in-domain RE10K data. In comparison, VD3D-DiT (the closest baseline) loses ${\approx}{20}$% of its visual fidelity on MSR-VTT.
@@
As Sec. 3.3 tells, we use the truncated normal distribution with location of 0.8 and scale of 0.075 with the [0.6, 1] bounds for training AC3D. We ablated the importance of biasing the noise sampling towards high noise and observe higher motion, visual quality, and camera controllability.
@@
We assess the importance of classifier-free guidance [44] on the camera conditioning in Tab. 2 (w/o camera CFG). It attains the same visual quality on both in-distribution (RE10K) and out-of-distribution (MSR-VTT) data, but degrades camera following, resulting in ${\approx}{5}$% higher pose reconstruction errors.
@@
To understand how well our curated data with scene motion but stationary cameras mitigate static scene bias, we train AC3D exclusively on RE10K, and report the results in Tab. 2 (w/o our dynamic data). It let the model maintain similar visual quality and text alignment on RE10K (in-domain data), but degraded its performance on out-of-distribution concepts of MSR-VTT (${\approx}{17}$% worse FID and ${\approx}{4}$% worse FVD). Evaluating the quality of scene motion is only reasonably through qualitative assessment, which is why we refer the reader to our qualitative comparisons in the supplementary which highlight the positive influence of our curated data on scene dynamics.
@@
We train AC3D using the original RE10K’s camera parameters without our scaling procedure and present the results in Tab. 2 (w/o metric scaled data). Since it is a more ambiguous conditioning signal, it made the model lose both visual quality (${\approx}{10}$% FVD on RE10K) and also fail to accurately follow the camera, leading to ${\approx}{12}$% higher trajectory reconstruction errors.
@@
As discussed in Sec. 3.6, we chose not to input the context information (text embeddings, resolution conditioning, etc.) into the lightweight camera branch to avoid potential interference with the camera representations. As Tab. 2 (w/o dropping camera context) shows, providing this information indeed results in ${\approx}{4}$% worse camera following and ${\approx}{15}$% lower visual quality.
@@
To mitigate visual quality and scene motion degradation, we attempted to perform joint training on 2D video data (without camera annotations) which was used in base VDiT training by applying dropout on camera inputs for it. While some prior works showed beneficial performance of such strategy [150] and, as Tab. 2 (with 2D training) shows, it indeed helped to maintain slightly higher visual fidelity (${\approx}{3}$% better FVD on MSR-VTT), camera steering severely deteriorated: up to 3×3$\times$3 × worse results for translation/rotation errors.
@@ 5 Conclusions
Our findings demonstrate that principled analysis of camera motion in video diffusion models leads to significant improvements in control precision and efficiency. Through enhanced conditioning schedules, targeted layer-specific camera control, and better-calibrated training data, AC3D achieves state-of-the-art performance in 3D camera-controlled video synthesis while maintaining high visual quality and natural scene dynamics. This work establishes a foundation for more precise and efficient camera control in text-to-video generation. We discuss the limitations of our approach in Appendix B. In future work, we plan to focus on further improving data limitations and develop control mechanisms for camera trajectories far outside of the training distribution.
@@ Appendix B Limitations
As discussed in Appendix D, we estimate the optical flows in the latent space of CogVideoX [167] autoencoder rather than the pixel space, because it’s the space the video DiT operates in and early trajectory steps produce disarranged decoder’s outputs. Besides (as also observed by [75]), motion spectral volumes are sensitive to the quality of optical flow estimation. While the key conclusions (of the camera motion being low-frequency and kicking in very early in the diffusion trajectory) would hold since the are evident even with a bare eye from inspecting the denoising process visualization, the exact behavioral details of motion spectral volumes might change when an optical flow estimator is swapped or the analysis is moved from the latent to pixel space.
@@
Appendix C Implementation details
@@
C.1 VDiT implementation details
@@
C.2 VDiT-CC implementation details
@@
C.3 AC3D implementation details
We train our complete setup for 6K iterations on the joint dataset of 65K videos from RealEstate10K [197] and 20K dynamic videos with static cameras (this dataset is described in Section 3.5 and Appendix F). We train with the learning rate of 0.0001 using the AdamW [89] optimizer with weight decay of 0.01 and cosine learning schedule [90].
@@ C.3 AC3D implementation details
Appendix D Motion analysis details

### v4

@@ 3.5 Mitigating training data limitations
Estimating camera parameters from in-the-wild videos remains challenging, as leading methods like [194, 117, 116, 147] frequently fail when processing videos containing dynamic scene content. This limitation results in camera-annotated datasets being heavily biased toward static scenes, which is particularly evident in RealEstate10K (RE10K) [201], the predominant dataset for training camera-controlled video models [151, 41, 6]. We hypothesize that models fine-tuned on such data interpret camera position information as a signal to suppress scene dynamics. This bias persists even when jointly training on unconstrained 2D video data [152], because the camera conditioning branch is only activated when camera parameters are available, which occurs exclusively for static scenes from RE10K, as static scenes remain the only reliable source for accurate camera annotation.
To address this fundamental limitation, we propose an alternative approach: rather than attempting to annotate dynamic scenes, which proved unsuccessful in our extensive preliminary research, even with state-of-the-art methods [147], we curate a collection of $20$K diverse videos featuring dynamic scenes captured by stationary cameras (see Figure 6). With stationary cameras, the camera position is inherently known, (we can assign fixed arbitrary extrinsics), allowing us to maintain active camera conditioning during training. This approach enables the camera conditioning branch to remain active during training while exposing the model to dynamic content, helping it distinguish between viewpoint conditioning and scene stillness. On top of this secondary dataset, following [152], we remove the scale ambiguity in RE10K by leveraging an off-the-shelf metric depth estimator; see Appendix H. Our experiments in Sec. 4.3 demonstrate that this straightforward yet effective data curation strategy successfully mitigates the distributional limitations of RE10K, restoring much of the lost scene dynamics, while maintaining precise camera control.
@@
Figure 6: RealEstate10k [201] videos (upper 2 rows) contain diverse camera trajectories, but are strongly biased towards static scenes. To mitigate this bias and also increase the concepts diversity, we curate $20$K videos with stationary cameras, but dynamic content (lower 2 rows). Such datasets are easy to construct, and surprisingly effective. Section 4.3 shows that integrating the dataset into our training improves visual quality on out-of-distribution prompts by $17\%$.
@@ 3.6 Miscellaneous improvements
$\hat{s}(\bm{x}|\bm{t},\bm{c})$ $=(1+w_{y}+w_{c})s_{\theta}(\bm{x}|\bm{t},\bm{c})$ (1)
$-w_{y}s_{\theta}(\bm{x}|\bm{c})-w_{c}s_{\theta}(\bm{x}|\bm{t}),$
where $\hat{s}(.)$ denotes the final update direction used during synthesis, $s_{\theta}$ represents the model’s predicted update direction, $\bm{t}$ and $\bm{c}$ are text and camera conditions, and $w_{y}$ and $w_{c}$ are their respective CFG weights. We zero-out the tensor for unconditional generation.
@@
Traditional ControlNet [190] conditioning, used in recent camera control methods [151, 41, 6], only processes conditioning signals without accessing the main branch. Our experiments show that using a bidirectional ControlNet produces better camera representations. This modification behaves as a feedback mechanism [73] provided by the main synthesis branch to the camera processing branch.
@@
Applying cross-attention over the context information (text prompts, resolution, etc.) in the camera DiT-XS blocks worsens visual quality and camera steering due to harmful interference of the context embeddings with camera representations.
@@ 4 Experiments
Our base VDiT model was trained on a large-scale dataset of text-annotated images and videos. VDiT-CC is fine-tuned from VDiT on RealEstate10K [201], contains ${\approx}{65}$M video clips with per-frame camera parameters since it is the setup used by existing methods [151, 41, 6].
@@
To assess the performance, we rely on a wide range of automatic quantitative metrics. We use FID [45], FVD [138], and CLIP score [44] to evaluate visual quality, and rotation and normalized translation errors [41] of ParticleSfM [194]-reconstructed trajectories to assess camera steerability. We evaluate them both on RE10K and MSR-VTT [163], since the latter allows to assess zero-shot generalization on out-of-distribution data. Moreover, we conduct a user study with details in the appendix in Sec. J.
@@ 4.1 Baselines
We select three camera-control methods: MotionCtrl [151], CameraCtrl [41], and VD3D [6]. MotionCtrl and CameraCtrl use a UNet-based video diffusion backbone [39], while VD3D builds on top of FIT [95, 23] and as such, is easily extendable to our video DiT [102] setup. Hence, we re-implement VD3D on top of our VDiT model to obtain an additional “VD3D+DiT” baseline. Moreover, we provide comparisons for an open-source model, i.e., CogVideoX [171]. See Sec. C of the appendix for more details.
Table 2: Quantitative evaluation. We evaluate all the models using camera pose and visual quality metrics based on unseen camera trajectories. We compute translation and rotation errors based on the estimated camera poses from generations using ParticleSfM [194]. We evaluate both in-distribution performance with RealEstate10K [201] and out-of-distribtion performance with MSR-VTT [163].
@@
MotionCtrl (U-Net) 0.477 0.094 2.99 61.70 26.46 0.593 0.137 16.85 283.12 24.11
CameraCtrl (U-Net) 0.465 0.089 2.48 55.64 26.81 0.587 0.132 12.33 201.33 25.05
VD3D (FIT) 0.409 0.043 1.40 42.43 28.07 0.504 0.050 7.80 165.18 26.89
VD3D (CogVideoX) 0.467 0.063 1.66 43.14 28.08 0.501 0.068 7.45 148.11 27.65
AC3D (CogVideoX) (ours) 0.374 0.039 1.27 38.20 28.62 0.431 0.039 5.52 116.04 28.38
MotionCtrl (VDiT) 0.504 0.126 1.74 43.81 27.69 0.589 0.146 9.92 150.20 27.25
CameraCtrl (VDiT) 0.513 0.138 1.62 42.10 27.73 0.566 0.143 8.15 146.77 27.51
VD3D (VDiT) 0.421 0.056 1.21 38.57 28.34 0.486 0.047 6.88 137.62 27.90
AC3D (VDiT) (ours) 0.358 0.035 1.18 36.55 28.76 0.428 0.038 5.34 110.71 28.58
@@ 4.2 Main results
We present quantitative comparisons with the baselines in Tab. 2. One can observe that just switching from the 4B-parameter pixel-space FIT [23] backbone, employed by the original VD3D approach, to our larger 11.5B-parameter latent-space DiT yields clear improvements across most metrics. Next, the results demonstrate that AC3D establishes a new state-of-the-art in performance against all baselines. Evaluating the quality of camera motion from still images is difficult, so we instead visualize all qualitative results in the website provided within our supplementary material. Therein, we can observe that AC3D better follows pose conditioning and achieves higher visual fidelity. We conduct user studies against VD3D+FIT (the original model) and VD3D+DiT (our improved re-implementation on top of the bigger video transformer). The results are presented in Table 1: AC3D outperforms them across all qualitative aspects, achieving a 90%+ overall preference score. Finally, we encourage the reader to assess the visual quality by observing videos on our website.
@@ 4.3 Ablations
The first ablation we conduct is to drop all camera conditioning, which makes the model equivalent to the vanilla VDiT. This is needed to understand the degradation of visual quality and text alignment. The results (Tab. 2, row w/o camera cond) show that our model loses less only ${\approx}{7}$% of the original visual fidelity on MSR-VTT (as measured by FVD), while (as expected) greatly improving on its in-domain RE10K data. In comparison, VD3D-DiT (the closest baseline) loses ${\approx}{20}$% of its visual fidelity on MSR-VTT.
@@
As Sec. 3.3 shows, we use the truncated normal distribution with location of 0.8 and scale of 0.075 with the [0.6, 1] bounds for training AC3D. We ablate the importance of biasing the noise sampling towards high noise and observe higher motion, visual quality, and camera controllability.
@@
We assess the importance of classifier-free guidance [46] on the camera conditioning in Tab. 2 (w/o camera CFG). It attains the same visual quality on both in-distribution (RE10K) and out-of-distribution (MSR-VTT) data, but degrades camera following, resulting in ${\approx}{5}$% worse pose reconstruction errors.
@@
To understand how well our curated data with scene motion but stationary cameras mitigates static scene bias, we train AC3D exclusively on RE10K, and report the results in Tab. 2 (w/o our dynamic data). The model maintains similar visual quality and text alignment on RE10K (in-domain data), but performance on out-of-distribution samples from MSR-VTT worsens (${\approx}{17}$% worse FID and ${\approx}{4}$% worse FVD). The quality of scene motion is better assessed by referring to our qualitative video comparisons in the supplementary.
@@
We train AC3D using the original RE10K’s camera parameters without our scaling procedure and present the results in Tab. 2 (w/o metric scaled data). This is a more ambiguous conditioning signal, and results in worse visual quality (${\approx}{10}$% FVD on RE10K) and camera following performance (${\approx}{12}$% worse trajectory reconstruction).
@@
As discussed in Sec. 3.6, we chose not to input the context information (text embeddings, resolution conditioning, etc.) into the camera branch to avoid potential interference with the camera representations. As Tab. 2 (w/o dropping camera context) shows, providing this information indeed results in ${\approx}{4}$% worse camera following and ${\approx}{15}$% lower visual quality.
@@
To mitigate visual quality and scene motion degradation, we attempted to perform joint fine-tuning on 2D video data (without camera annotations) which was used in base VDiT training by applying dropout on camera inputs for it. Prior work shows performance benefits with this strategy [152] and, as Tab. 2 (with 2D training) shows, it indeed helps to maintain slightly higher visual fidelity in our case (${\approx}{3}$% better FVD on MSR-VTT). However, camera steering severely deteriorates, leading to up to 3×3$\times$3 × worse results for translation/rotation errors.
@@ 5 Conclusions
Our findings demonstrate that principled analysis of camera motion in video diffusion models leads to significant improvements in control precision and efficiency. Through enhanced conditioning schedules, targeted layer-specific camera control, and better-calibrated training data, AC3D achieves state-of-the-art performance in 3D camera-controlled video synthesis while maintaining high visual quality and natural scene dynamics. This work establishes a foundation for more precise and efficient camera control in text-to-video generation. We discuss the limitations of our approach in Appendix B. In future work, we plan to focus on further improving data limitations and developing control mechanisms for camera trajectories far outside of the training distribution.
6 Acknowledgements
DBL acknowledges support from NSERC under the RGPIN program, the Canada Foundation for Innovation, and the Ontario Research Fund.
@@ Appendix B Limitations
As discussed in Appendix E, we estimate the optical flows in the latent space of CogVideoX [170] autoencoder rather than the pixel space, because it’s the space the video DiT operates in and early trajectory steps produce disarranged decoder’s outputs. Besides (as also observed by [77]), motion spectral volumes are sensitive to the quality of optical flow estimation. While the key conclusions (of the camera motion being low-frequency and kicking in very early in the diffusion trajectory) would hold since the are evident even with a bare eye from inspecting the denoising process visualization, the exact behavioral details of motion spectral volumes might change when an optical flow estimator is swapped or the analysis is moved from the latent to pixel space.
@@
Appendix C CogVideoX Results
Figure 7: Comparing the average magnitude of motion spectral volumes for scenes with different motion types for CogVideoX [171]. Videos with camera motion (purple) exhibit stronger overall motion than the videos with scene motion (orange), especially for the low-frequency range, suggesting that the motion induced by camera transitions is heavily biased towards low-frequency components.
Figure 8: How camera motion is modeled by diffusion (CogVideoX)? As visualized in Figure 4(a) and Figure 3, the motion induced by camera transitions is a low-frequency type of motion. We observe that a video DiT creates low-frequency motion very early in the denoising trajectory: Figure 4(b) (left) shows that even at $t{=}0.96$ (first ${\approx}{4}$% of the steps), the low-frequency motion components have already been created, while high frequency ones do not fully unveil even till $t{=}0.5$. We found that controlling the camera pose later in the denoising trajectory is not only unnecessary but detrimental to both scene motion and overall visual quality. Here, we provide the same analysis conducted for CogVideox [171].
We implement our method on top of CogVideoX [171] to show generalizability. Moreover, we conduct the motion analysis of the main paper for CogVideoX and show results in Fig. 7 and Fig. 8. We observe a similar pattern, confirming the generalizability of our findings.
Appendix D Implementation details
@@
D.1 VDiT implementation details
@@
D.2 VDiT-CC implementation details
@@
D.3 AC3D implementation details
We train our complete setup for 6K iterations on the joint dataset of 65K videos from RealEstate10K [201] and 20K dynamic videos with static cameras (this dataset is described in Section 3.5 and Appendix G). We train with the learning rate of 0.0001 using the AdamW [91] optimizer with weight decay of 0.01 and cosine learning schedule [92].
@@ D.3 AC3D implementation details
1D temporal camera encoder
VDiT-CCprocesses videos in the autoencoder’s latent space with 4×4$\times$4 × temporal compression [170], raising the question of how to incorporate full-resolution camera parameters. While camera poses could be naively downsized (e.g., subsampled or reshaped) to match the latent resolution, this would force small DiT-XS blocks to process compressed camera information. Instead, we implement a sequence of causal 1D convolutions that transform a $F{\times}6$ sequence of Plücker coordinates for each pixel into a $(F//4){\times}32$ representation.
Appendix E Motion analysis details